OpenStudy (anonymous):

3x^2 + 41x +k = 0 the product of roots is 7

OpenStudy (alekos):

product of the roots equals c/a

OpenStudy (gorv):

ax^2+bx+c =0 this is standard equation in this product of roots =c/a

OpenStudy (gorv):

in your equation a=3 and c= k

OpenStudy (gorv):

and product=c/a=7 given so put values of c and a and solve